Find an equation of the line that passes through the point (5, 3) and is perpendicular to the line
4x + 9y - 1 = 0

Answer by TimothyLamb(4379)

---
given:
4x + 9y - 1 = 0
---
find perpendicular slope:
---
9y = -4x + 1
y = (-4/9)x + 1/9
---
perpendicular slope = 9/4
---
find line:
through: (5, 3)
slope: 9/4 = 2.25
---
to solve the linear equation, copy and paste this (the given point and perpendicular slope):
5, 3, 2.25
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (3.66666667, 0)
y-intercept = (0, -8.25)
slope = 9/4 = 2.25
---
answer:
the slope-intercept form of our new perpendicular line:
y = (9/4)x - 8.25
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php